# Grounding DINO: Marrying DINO with Grounded Pre-Training for Open-Set Object Detection

### v3

@@ 4 Experiments
4.1 Setup
We conduct extensive experiments on three settings: a closed-set setting on the COCO detection benchmark (Sec. C.1), an open-set setting on zero-shot COCO, LVIS, and ODinW (Sec. 4.2), and a referring detection setting on RefCOCO/+/g (Sec. 4.3). Ablations are then conducted to show the effectiveness of our model design (Sec. 4.4). We also explore a way to transfer a well-trained DINO to the open-set scenario by training a few plug-in modules in Sec. 4.5. The test of our model efficiency is presented in Sec. I.
Implementation Details We trained two model variants, Grounding-DINO-T with Swin-T [32], and Grounding-DINO-L with Swin-L [32] as an image backbone, respectively. We leveraged BERT-base [8] from Hugging Face [51] as text backbones. As we focus more on the model performance on novel classes, we list zero-shot transfer and referring detection results in the main text. More implementation details are available in the Appendix Sec. A.
@@ 4.2 Zero-Shot Transfer of Grounding DINO
We compare Grounding DINO with GLIP and DINO in Table 2. We pre-train models on large-scale datasets and directly evaluate our model on the COCO benchmark. As the O365 dataset [44] has (nearly [^4] It is not an exact mapping between O365 and COCO categories. We made some approximations during evaluation. ) covered all categories in COCO, we evaluate an O365 pre-trined DINO on COCO as a zero-shot baseline. The result shows that DINO performs better on the COCO zero-shot transfer than DyHead. Grounding DINO outperforms all previous models on the zero-shot transfer setting, with +0.5+0.5AP and +1.8+1.8AP compared with DINO and GLIP under the same setting. Grounding data is still helpful for Grounding DINO, introducing more than 11AP (48.1 vs. 46.7) on the zero-shot transfer setting. With stronger backbones and larger data, Grounding DINO sets a new record of $52.5$ AP on the COCO object detection benchmark without seeing any COCO images during training. Grounding DINO obtains a $62.6$ AP on COCO minival, outperforming DINO’s $62.5$ AP. When enlarging the input images by $1.5\times$, the benefits reduce. We suspect that the text branch enlarges the gap between models with different input images. Even though, Grounding DINO gets an impressive $63.0$ AP on COCO test-dev with fine-tuning on the COCO dataset(See the number in brackets of Table 2).
@@
LVIS [14] is a dataset for long-tail objects. It contains more than $1000$ categories for evaluation. We use LVIS as a downstream task to test the zero-shot abilities of our model. We use GLIP as baselines for our models. The results are shown in Table 3. Grounding DINO outperforms GLIP under the same settings. We found two interesting phenomena in the results. First, Grounding DINO works better than common objects than GLIP, but worse on rare categories. We suspect that the $900$ query design limits the ability for long-tailed objects. By contrast, the one-stage detector uses all proposals in the feature map for comparisons. The other phenomenon is that Grounding DINO has larger gains with more data than GLIP. For example, Grounding DINO introduces +1.8+1.8 AP gains with the caption data Cap4M, whereas GLIP has only +1.1+1.1 AP. We believe that Grounding DINO has a better scalability compared with GLIP. A larger-scale training will be left as our future work.
Table 4: Results on the ODinW benchmark.
@@
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 20.0 9.5
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG,Cap4M 22.3 11.9
@@
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 46.4 51.1
@@
Grounding-DINO-T(Ours) Swin-T 172M O365,GoldG 70.7 76.2
@@
ODinW (Object Detection in the Wild) [23] is a more challenging benchmark to test model performance under real-world scenarios. It collects more than $35$ datasets for evaluation. We report three settings, zero-shot, few-shot, and full-shot results in Table 4. Grounding DINO performs well on this benchmark. With only O365 and GoldG for pre-train, Grounding-DINO-T outperforms DINO on few-shot and full-shot settings. Impressively, Grounding DINO with a Swin-T backbone outperforms DINO with Swin-L on the full-shot setting. Grounding DINO outperforms GLIP under the same backbone for the zero-shot setting, comparable with GLIPv2 [59] without any new techniques like masked training. The results show the superiority of our proposed models. Grounding-DINO-L set a new record on ODinW zero-shot with a $26.1$ AP, even outperforming the giant Florence models [55]. The results show the generalization and scalability of Grounding DINO.
@@
Grounding-DINO-T (Ours) Swin-T O365,GoldG 50.41 57.24 43.21 51.40 57.59 45.81 67.46 67.13
Grounding-DINO-T (Ours) Swin-T O365,GoldG,RefC 73.98 74.88 59.29 66.81 69.91 56.09 71.06 72.07
Grounding-DINO-T (Ours) Swin-T O365,GoldG,RefC 89.19 91.86 85.99 81.09 87.40 74.71 84.15 84.94
Grounding-DINO-L (Ours)* Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 90.56 93.19 88.24 82.75 88.95 75.92 86.13 87.02
@@
Model Pre-Train Data COCO minival LVIS minival ODinW
DINO Pre-Train Grounded Fine-Tune Zero-Shot Zero-Shot Zero-Shot
Grounding-DINO-T - O365 46.7 16.2 14.5
(from scratch) - O365,GoldG 48.1 25.6 20.0
Grounding-DINO-T O365 O365 46.5 17.9 13.6
(from pre-trained DINO) O365 O365,GoldG 46.4 26.1 18.5
4.3 Referring Object Detection Settings
4.4 Ablations
We conduct ablation studies in this section. We propose a tight fusion grounding model for open-set object detection and a sub-sentence level text prompt. To verify the effectiveness of the model design, we remove some fusion blocks for different variants. Results are shown in Table 6. All models are pre-trained on O365 with a Swin-T backbone. The results show that each fusion helps the final performance. Encoder fusion is the most important design. The impact of word-level text prompts the smallest, but helpful as well. The language-guided query selection and text cross-attention present a larger influence on LVIS and COCO, respectively.
4.5 Transfer from DINO to Grounding DINO
Recent work has presented many large-scale image models for detection with DINO architecture [^6] See model instances at https://github.com/IDEA-Research/detrex . It is computationally expensive to train a Grounding DINO model from scratch. However, the cost can be significantly reduced if we leverage pre-trained DINO weights. Hence, we conduct some experiments to transfer pre-trained DINO to Grounding DINO models. We freeze the modules co-existing in DINO and Grounding DINO and fine-tune the other parameters only. (We compare DINO and Grounding DINO in Sec. E.) The results are available in Table 7.
It shows that we can achieve similar performances with Grounding-DINO-Training only text and fusion blocks using a pre-trained DINO. Interestingly, the DINO-pre-trained Grounding DINO outperforms standard Grounding DINO on LVIS under the same setting. The results show that there might be much room for model training improvement, which will be our future work to explore. With a pre-trained DINO initialization, the model converges faster than Grounding DINO from scratch, as shown in Fig. 5. Notably, we use the results without exponential moving average (EMA) for the curves in Fig. 5, which results in a different final performance that in Table 7. As the model trained from scratch need more training time, we only show results of early epochs.
Figure 5: Comparison between two Grounding DINO variants: Training from scratch and transfer from DINO-pretrained models. The models are trained on O365 and evaluated on COCO directly.
@@ 5 Conclusion
Limitations: Although the great performance on open-set object detection setting, Grounding DINO cannot be used for segmentation tasks like GLIPv2. Moreover, our training data is less than the largest GLIP model, which may limit our final performance.
6 Acknowledgement
We thank the authors of GLIP [26]: Liunian Harold Li, Pengchuan Zhang, and Haotian Zhang for their helpful discussions and instructions. We also thank Tiancheng Zhao, the author of OmDet [61], and Jianhua Han, the author of DetCLIP [53], for their response on their model details. We thank He Cao of The Hong Kong University of Science and Technology for his helps on diffusion models.
Appendix A More Implementation Details
[…]
@@ Appendix A More Implementation Details
Appendix B Data Usage
@@
    Caption data. To enhance the model performance on novel categories, we feed the semantic-rich caption data to our model. Following GLIP, we use the pseudo-labeled caption data for model training. A well-trained model generates the pseudo labels.
There are two versions of the O365 dataset, which we termed O365v1 and O365v2, respectively. O365v1 is a subset of O365v2. O365v1 contains about 600K images, while O365v2 contains about 1.7M images. Following previous works [26, 53], we pre-train the Grounding-DINO-T on O365v1 for a fair comparison. The Grounding-DINO-L is pre-trained on O365v2 for a better result.
Appendix C More Results on COCO Detection Benchmarks
@@
Appendix D Detailed Results on ODinW
We present detailed results of Grounding DINO on ODinW35 in Table 10, Table 11, and Table 12.
@@ Appendix D Detailed Results on ODinW
Appendix E Comparison between DINO and Grounding DINO
To illustrate the difference between DINO and Grounding DINO, we compare DINO and Grounding DINO in Fig. 6. We mark the DINO blocks in gray, while the newly proposed modules are shaded in blue.

### v5

@@ 4 Experiments
We conduct extensive experiments on three settings: a closed-set setting on the COCO detection benchmark (Sec. 0.C.2), an open-set setting on zero-shot COCO, LVIS, and ODinW (Sec. 4.2), and a referring detection setting on RefCOCO/+/g (Sec. 4.3). Ablations are then conducted to show the effectiveness of our model design (Sec. 4.5). We also explore a way to transfer a well-trained DINO to the open-set scenario by training a few plug-in modules in Sec. 0.C.1. The test of our model efficiency is presented in Sec. 0.C.4.
4.1 Implementation Details
We trained two model variants, Grounding DINO T with Swin-T [32], and Grounding DINO L with Swin-L [32] as an image backbone, respectively. We leveraged BERT-base [8] from Hugging Face [49] as text backbones. As we focus more on the model performance on novel classes, we list zero-shot transfer and referring detection results in the main text.
[…]
@@ 4.2 Zero-Shot Transfer of Grounding DINO
We compare Grounding DINO with GLIP and DINO in Table 2. We pre-train models on large-scale datasets and directly evaluate our model on the COCO benchmark. As the O365 dataset [43] has (nearly [^4] It is not an exact mapping between O365 and COCO categories. We made some approximations during evaluation. ) covered all categories in COCO, we evaluate an O365 pre-trined DINO on COCO as a zero-shot baseline. The result shows that DINO performs better on the COCO zero-shot transfer than DyHead. Grounding DINO outperforms all previous models on the zero-shot transfer setting, with +0.5+0.5AP and +1.8+1.8AP compared with DINO and GLIP under the same setting. Grounding data is still helpful for Grounding DINO, introducing more than 11AP (48.1 vs. 46.7) on the zero-shot transfer setting. With stronger backbones and larger data, Grounding DINO sets a new record of $52.5$ AP on the COCO object detection benchmark without seeing any COCO images during training. Grounding DINO obtains a $62.6$ AP on COCO minival, outperforming DINO’s $62.5$ AP. When enlarging the input images by $1.5\times$, the benefits reduce. We suspect that the text branch enlarges the gap between models with different input images. Even though the performance plateaus with larger input size, Grounding DINO gets an impressive $63.0$ AP on COCO test-dev with fine-tuning on the COCO dataset(See the number in brackets of Table 2).
@@
LVIS [15] is a dataset for long-tail objects. It contains more than $1000$ categories for evaluation. We use LVIS as a downstream task to test the zero-shot abilities of our model. We use GLIP and DetCLIPv2 as baselines for our models. The results are shown in Table 3.
Model Backbone Pre-Training Data MiniVal [18]
AP APr/APc/APf
Zero-Shot Setting
GLIP-T (C) Swin-T O365,GoldG 24.9 17.7/19.5/31.0
GLIP-T Swin-T O365,GoldG,Cap4M 26.0 20.8/21.4/31.0
DetCLIPv2 Swin-T O365,GoldG,CC15M 40.4 36.0/41.7/40.0
Grounding DINO T Swin-T O365,GoldG 25.6 14.4/19.6/32.2
Grounding DINO T Swin-T O365,GoldG,Cap4M 27.4 18.1/23.3/32.7
Grounding DINO L Swin-L O365,OI,GoldG,Cap4M, COCO,RefC 33.9 22.2/30.7/38.8
Fine-Tune Setting
MDETR RN101 GoldG,RefC 24.2 20.9/24.9/24.3
Mask R-CNN RN101 - 33.3 26.3/34.0/33.9
DetCLIPv2 [51] Swin-T O365,GoldG,CC15M 50.7 44.3/52.4/50.3
Grounding DINO T Swin-T O365,GoldG 52.1 35.4/51.3/55.7
Table 3: Model results on LVIS.
We found two interesting phenomena in the results. First, Grounding DINO works better than common objects than GLIP, but worse on rare categories. We reviewed DETR-like models on LVIS and noted these models often exhibit lower rare category AP despite similar overall AP, like Table 2 of [9] and Table 6 of [18]. To our knowledge, no existing DETR-like models effectively address the rarity challenge in LVIS without extra training data, which may be a characteristic limitation of the architecture.
The other phenomenon is that Grounding DINO has larger gains with more data than GLIP. For example, Grounding DINO introduces +1.8+1.8 AP gains with the caption data Cap4M, whereas GLIP has only +1.1+1.1 AP. We believe that Grounding DINO has better scalability compared with GLIP. A larger-scale training will be left as our future work.
Although achieving better results than GLIP, we found that Grounding DINO is inferior to DetCLIPv2, which is trained on a larger scale data. This performance difference might be attributed to the disparity in data distribution between the training dataset and the LVIS dataset.
To unveil the full potential of Grounding DINO, we fine-tuned it on the LVIS dataset. Table 3 highlights the commendable capability of our model. Remarkably, despite being pre-trained only on the O365 and GoldG datasets, Grounding DINO outperforms DetCLIPv2-T by a margin of $1.5$ AP. This result shows that Grounding DINO might have learned a better object-level representation which helps yield a better performance after fine-tuning (aligning with the target dataset). In our future work, we will perform more studies, including varying the semantic concept coverage of the training data and increasing the scale of the training data, to further improve the zero-shot generalization performance.
Table 4: Model results on the ODinW benchmark.
@@
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 20.0 9.5
Grounding DINO T(Ours) Swin-T 172M O365,GoldG,Cap4M 22.3 11.9
@@
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 46.4 51.1
@@
Grounding DINO T(Ours) Swin-T 172M O365,GoldG 70.7 76.2
@@
ODinW (Object Detection in the Wild) [23] is a more challenging benchmark to test model performance under real-world scenarios. It collects more than $35$ datasets for evaluation. We report three settings, zero-shot, few-shot, and full-shot results in Table 4. Grounding DINO performs well on this benchmark. With only O365 and GoldG for pre-train, Grounding DINO T outperforms DINO on few-shot and full-shot settings. Impressively, Grounding DINO with a Swin-T backbone outperforms DINO with Swin-L on the full-shot setting.
Grounding DINO outperforms GLIP under the same backbone for the zero-shot setting. Grounding DINO and GLIPv2-T show similar $AP_{average}$. However, a key distinction lies in the $AP_{median}$, where Grounding DINO significantly outperforms GLIPv2-T (11.9 vs 8.9). This suggests that while GLIPv2 may exhibit larger performance variance across different datasets, Grounding DINO maintains a more consistent performance level. GLIPv2 incorporates advanced techniques like masked text training and cross-instance contrastive learning, making it more complex than our Grounding DINO model. Moreover, our model is more compact (172M parameters) compared to GLIPv2 (232M parameters). These factors combined—performance consistency, model complexity, and size—should address concerns about our model’s capability in true open-set scenarios.
Grounding DINO L set a new record on ODinW zero-shot with a $26.1$ AP, even outperforming the giant Florence models [54]. The results show the generalization and scalability of Grounding DINO.
4.3 Referring Object Detection Settings
@@
Grounding DINO T (Ours) Swin-T O365,GoldG 50.41 57.24 43.21 51.40 57.59 45.81 67.46 67.13
Grounding DINO T (Ours) Swin-T O365,GoldG,RefC 73.98 74.88 59.29 66.81 69.91 56.09 71.06 72.07
Grounding DINO T (Ours) Swin-T O365,GoldG,RefC 89.19 91.86 85.99 81.09 87.40 74.71 84.15 84.94
Grounding DINO L (Ours)* Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 90.56 93.19 88.24 82.75 88.95 75.92 86.13 87.02
@@
4.4 Effects of RefC and COCO Data
We add the RefCOCO/+/g (we note it as “RefC” in tables) and COCO into training in some settings. We explore the influence of these data in Table 6. The results show that RefC helps improve the COCO zero-shot and fine-tuning performance but hurts the LVIS and ODinW results. With COCO introduced, the COCO results are greatly improved. It shows that COCO brings marginal improvements in LVIS and slight decreases on ODinW.
Model Pre-Train COCO minival LVIS minival ODinW
Zero-Shot Fine-Tune Zero-Shot Zero-Shot
Grounding DINO T O365,GoldG 48.1 57.1 25.6 20.0
Grounding DINO T O365,GoldG,RefC 48.5 57.3 21.9 17.7
Grounding DINO T O365,GoldG,RefC,COCO 56.1 57.5 22.3 17.4
4.5 Ablations
We conduct ablation studies in this section. We propose a tight fusion grounding model for open-set object detection and a sub-sentence level text prompt. To verify the effectiveness of the model design, we remove some fusion blocks for different variants. Results are shown in Table 7. All models are pre-trained on O365 with a Swin-T backbone.
The results show that encoder fusion significantly improves model performance on both COCO and LVIS datasets. The results from comparing model #\#1 with the baseline model #\#0 validate this observation. Other techniques, such as language-guided query selection, text cross-attention, and sub-sentence text prompt, also contribute positively to the LVIS performance, yielding significant gains of +3.0 AP, +1.8 AP, and +0.5 AP, respectively. Additionally, these methods enhance the COCO zero-shot performance, further underscoring their effectiveness. However, we observed that language-guided query selection and sub-sentence text prompt had minimal impact on the COCO fine-tune performance. This outcome is reasonable, given that these methods do not alter model parameters or add computational burdens. Text cross-attention, while introducing fewer parameters than encoder fusion, showed less performance improvement compared to encoder fusion (+0.6 vs. +0.8). This finding suggests that fine-tuning performance is predominantly influenced by the model’s parameters, indicating that scaling models is a promising direction for enhancing performance.
@@ 5 Conclusion
Limitations: Despite the great performance on open-set object detection settings, Grounding DINO cannot be used for segmentation tasks like GLIPv2. Our training data is less than the largest GLIP model, which may limit our final performance. Moreover, we find that our model will produce false positive results in some cases, which may need more techniques or data to reduce the hallucination.
Social Impacts: The use of deep learning models, such as this one, exposes them to vulnerabilities through adversarial attacks. Additionally, the accuracy and correctness of the model’s outputs cannot be guaranteed. There is also the risk that the open-set detection capabilities of the model could be exploited for unlawful purposes.
Acknowledgement
We thank the authors of GLIP [25]: Liunian Harold Li, Pengchuan Zhang, and Haotian Zhang for their helpful discussions and instructions. We also thank Tiancheng Zhao, the author of OmDet [60], and Jianhua Han, the author of DetCLIP [52], for their response to their model details. We thank He Cao of The Hong Kong University of Science and Technology for his help on diffusion models.
Appendix 0.A More Implementation Details
0.A.1 Hyperparameters
Table 8 presents the hyperparameters used in our main experiments.
@@ 0.A.1 Hyperparameters
0.A.2 Pseudo Code Language-Guided Query Selection
We present the pseudo-code of the Language-Guided Query Selection module in Algorithm 1.
1"""
2Input:
3image_feat:(bs,num_img_tokens,ndim)
4text_feat:(bs,num_text_tokens,ndim)
5num_query:int
6
7Output:
8topk_idx:(bs,num_query)
9"""
10
11logits = torch.einsum("bic,btc->bit", image_feat, text_feat) # bs, num_img_tokens, num_text_tokens
12logits_per_img_feat = logits.max(-1)[0]# bs, num_img_tokens
13topk_idx = torch.topk(logits_per_img_feature, num_query, dim=1)[1] # bs, num_query
Algorithm 1 Pseudocode of Language-guided Query Selection in PyTorch-like style.
The variables image_feat and text_feat are used for image and text features, respectively. num_query is the number of queries in the decoder, which is set to $900$ in our implementation. We use bs and ndim for batch size and feature dimension in the pseudo-code. num_img_tokens and num_text_tokens are used for the number of image and text tokens, respectively.
Appendix 0.B Data Usage
@@
    Caption data. To enhance the model performance on novel categories, we feed the semantic-rich caption data to our model. Following GLIP, we use the pseudo-labeled caption data for model training. In our experiments, we use the same data with GLIP under comparable settings. More specifically, we use GLIP-T annotated caption data for Grounding DINO T, while GLIP-L annotated caption data for Grounding DINO L.
There are two versions of the O365 dataset, which we termed O365v1 and O365v2, respectively. O365v1 is a subset of O365v2. O365v1 contains about 600K images, while O365v2 contains about 1.7M images. Following previous works [25, 52], we pre-train the Grounding DINO T on O365v1 for a fair comparison. The Grounding DINO L is pre-trained on O365v2 for a better result.
Appendix 0.C More Experiment Results
0.C.1 Transfer from DINO to Grounding DINO
Recent work has presented many large-scale image models for detection with DINO architecture [^6] See model instances at https://github.com/IDEA-Research/detrex . It is computationally expensive to train a Grounding DINO model from scratch. However, the cost can be significantly reduced if we leverage pre-trained DINO weights. Hence, we conduct some experiments to transfer pre-trained DINO to Grounding DINO models. We freeze the modules co-existing in DINO and Grounding DINO and fine-tune the other parameters only. (We compare DINO and Grounding DINO in Sec. 10.) The results are available in Table 9.
Model Pre-Train Data COCO minival LVIS minival ODinW
DINO Pre-Train Grounded Fine-Tune Zero-Shot Zero-Shot Zero-Shot
Grounding DINO T - O365 46.7 16.2 14.5
(from scratch) - O365,GoldG 48.1 25.6 20.0
Grounding DINO T O365 O365 46.5 17.9 13.6
(from pre-trained DINO) O365 O365,GoldG 46.4 26.1 18.5
It shows that we can achieve similar performances with Grounding DINO-Training only text and fusion blocks using a pre-trained DINO. Interestingly, the DINO-pre-trained Grounding DINO outperforms the standard Grounding DINO on LVIS under the same setting. The results show that there might be much room for model training improvement, which will be our future work to explore.
Table 10: Comparison between two Grounding DINO variants: Training from scratch and transfer from DINO-pretrained models. The models are trained on O365 and evaluated on COCO.
With a pre-trained DINO initialization, the model converges faster than Grounding DINO from scratch, as shown in Fig. 10. Notably, we use the results without exponential moving average (EMA) for the curves in Fig. 10, which results in a different final performance that in Table 9. As the model trained from scratch need more training time, we only show results of early epochs.
Comparison between DINO and Grounding DINO To illustrate the difference between DINO and Grounding DINO, we compare DINO and Grounding DINO in Fig. 5. We mark the DINO blocks in gray, while the newly proposed modules are shaded in blue.
0.C.2 Detailed Results on COCO Detection Benchmarks
@@
0.C.3 Detailed Results on ODinW Benchmarks
We present detailed results of Grounding DINO on ODinW35[23] in Table 12, Table 13, and Table 14.
@@ 0.C.3 Detailed Results on ODinW Benchmarks
0.C.4 Model Efficiency
We compare the model size and efficiency between Grounding DINO T and GLIP-T in Table 20. The results show that our model has a smaller parameter size and better efficiency than GLIP.
0.C.5 Ablations for More Decoder Queries
To verify the model performance with more decoder queries, we conducted additional experiments with 1200 and 1500 queries on the COCO and LVIS datasets, detailed in Table 15 below.
The results indicate that models with 1200 and 1500 queries slightly outperform the 900-query version on LVIS rare classes, suggesting better coverage of rare classes. However, the improvement is marginal, as the 900-query model already sufficiently covers all objects in both COCO and LVIS. Additionally, introducing more queries exacerbates data imbalance during training, as the model is trained on objects from sampled categories. This imbalance could offset the benefits of additional queries.
Pretrain Data Query Num COCO LVIS
AP APS APM APL AP APr APc APf
O365 900 46.7 32.1 50.0 61.3 15.8 9.4 22.9 28.4
O365 1200 46.7 32.3 49.9 61.1 15.7 9.0 23.1 29.1
O365 1500 46.9 32.7 50.3 61.3 15.8 9.2 23.0 29.0
Table 15: Results for Grounding DINO Tiny with more decoder queries
0.C.6 Results with Different Language Encoder for REC
To verify the impacts of language encoders with different sizes, we conducted experiments using two variants of BERT: bert-base-uncased (BERT-B) and bert-large-uncased (BERT-L). These models were trained on a combined dataset consisting of RefCOCO, RefCOCO+, and RefCOCOg. We removed the leaked data in the combined dataset for a fair comparison. It’s important to note that training on this combined dataset, as opposed to tuning each dataset separately, might result in slightly lower performance. For a fair comparison, we initialized the models with the O365+GoldG+Cap4M checkpoint, except for the BERT parameters. Limited by time and resources, the training duration was capped at 18 epochs.
Interestingly, our results showed that Grounding DINO with BERT-B outperformed or matched the BERT-L variant in most metrics (as shown in Table 16). This suggests that our default use of BERT-B during the pretrain stage may have contributed to its better performance. Moreover, we didn’t observe significant improvements in the late stages of training, indicating that both models were nearing their optimal performance.
This outcome suggests that the main limitation in enhancing REC performance lies within the detection branch, rather than the language processing module. A dedicated model for REC data might be helpful for REC tasks.
Model RefCOCO RefCOCO+ RefCOCOg
val testA testB val testA testB val test
Grounding DINO T (BERT-B) 87.4 91.6 84.2 78.6 86.5 73.4 81.6 83.3
Grounding DINO T (BERT-L) 87.0 91.4 84.0 78.1 86.4 73.0 81.7 83.3
Table 16: Results for Grounding DINO with different language encoders
0.C.7 Detic Pseudo-labeled Data for LVIS
To illustrate the potential of our model under similar conditions, we conducted oracle experiments as shown in Table 17. We pseudo-labeled the ImageNet dataset using a pre-trained Detic[62] model and filtered out LVIS-related images for training, creating a new pseudo-labeled dataset named IN22K-LVIS-1M. It contains about 1M pseudo-labeled images for training. Note that our model under the setting may be not a real zero-shot setting, as our pseudo labeler Detic is trained with LVIS data. The results from these experiments suggest that Grounding DINO can achieve promising LVIS performance, even without direct LVIS training data. A similarly distributed dataset can help the model to generalize well.
Model Pre-train Data LVIS MiniVal - AP(r/c/f)
Zero-shot Fine-tune
DetCLIPv2-T OG + CC15M 40.4 (36.0 / 41.7 / 40.0) 50.7 (44.3 / 52.4 / 50.3)
Grounding DINO T OG+IN22K-LVIS-1M 40.6 (38.5 / 41.1 / 40.4) 54.5 (47.3 / 53.9 / 56.1)
Table 17: Oracle experiments on LVIS. Note that our model under the setting may be not a real zero-shot setting, as our pseudo labeler Detic is trained with LVIS data.
0.C.8 Comparison between Grounding DINO and GLIP on ODinW
In our comparison of Grounding DINO and GLIP across various datasets in ODinW, as presented in Table 18, we observe that Grounding DINO underperforms on certain uncommon datasets where both models generally show limited effectiveness. For instance, in the PlantDoc dataset, Grounding DINO scores 0.36 compared to GLIP’s 1.1. This dataset includes infrequent categories such as "Tomato leaf mosaic virus," which are not well-represented in the training data. These findings highlight the need for improving data quality to enhance overall model performance.
Metric GLIP-T Grounding DINO T
Average Score (↑) 19.6 22.3
Median Score (↑) 5.1 11.9
AerialMaritimeDrone_large (↑) 13.70 10.30
AerialMaritimeDrone_tiled (↑) 12.60 17.50
AmericanSignLanguageLetters_American_Sign_Language_Letters (↑) 2.50 0.78
Aquarium_Aquarium_Combined (↑) 18.30 18.64
BCCD_BCCD (↑) 1.00 11.96
ChessPieces_Chess_Pieces (↑) 10.00 15.62
CottontailRabbits (↑) 69.70 67.61
DroneControl_Drone_Control (↑) 5.10 4.99
EgoHands_generic (↑) 50.00 57.64
EgoHands_specific (↑) 0.80 0.69
HardHatWorkers (↑) 3.00 4.05
MaskWearing (↑) 1.10 0.25
MountainDewCommercial (↑) 21.60 25.46
NorthAmericaMushrooms_North_American_Mushrooms (↑) 75.10 68.18
OxfordPets_by-breed (↑) 0.40 0.21
OxfordPets_by-species (↑) 1.10 1.30
PKLot_640 (↑) 0.00 0.06
Packages_Raw (↑) 72.30 60.53
PascalVOC (↑) 56.10 55.65
Raccoon_Raccoon (↑) 57.80 60.07
ShellfishOpenImages (↑) 25.90 29.56
ThermalCheetah (↑) 2.70 17.72
UnoCards (↑) 0.20 0.81
VehiclesOpenImages (↑) 56.00 58.49
WildfireSmoke (↑) 2.30 20.04
boggleBoards_416x416AutoOrient_export (↑) 0.00 0.29
brackishUnderwater (↑) 3.70 1.47
dice_mediumColor_export (↑) 1.10 0.33
openPoetryVision (↑) 0.00 0.05
pistols_export (↑) 49.80 66.99
plantdoc (↑) 1.10 0.36
pothole (↑) 17.20 25.21
selfdrivingCar_fixedLarge_export (↑) 8.00 9.95
thermalDogsAndPeople (↑) 43.70 67.89
websiteScreenshots (↑) 0.50 1.30
Table 18: Comparison of Grounding DINO and GLIP on ODinW. Both models are trained on O365, GoldG, and Cap4M with Swin-Tiny backbones.
0.C.9 Training Grounding DINO on RefCOCO/+/g from scratch
To demonstrate our model’s capabilities, we trained a Grounding DINO from scratch on the RefCOCO/+/g datasets, as shown in Table 19. Despite time limitations that restricted us to 9 epochs of training, we achieved comparable performance to SOTA REC models. This underscores the potential of our approach to handling REC tasks effectively.
refcoco (val/testA/testB) refcoco+ (val/testA/testB) refcocog (val/test)
TransVG R101 81.02/82.72/78.35 64.82/70.70/56.94 68.67/67.73
Grounding DINO T (9ep) SwinT 81.42/85.28/76.75 69.14/74.79/60.17 73.96/74.37
Table 19: Training on RefCOCO from scratch.
Appendix 0.D Visualizations
0.D.1 Detection Visualizations